OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (7, 4); Slope: 6 (Points : 1) y – 7 = 6(x + 4) y – 4 = 6(x – 7) y – 4 = 6(x + 7) y + 4 = 6(x – 7)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Here is our standard form for point-slope form of a line:\[\Large\rm y-\color{orangered}{y_1}=\color{green}{m}(x-\color{royalblue}{x_1})\] And they gave us this information:\[\Large\rm \color{green}{slope=6}\]\[\Large\rm (\color{royalblue}{7},\color{orangered}{4})\]

zepdrix (zepdrix):

Do you understand how to plug the information in? Pssst! Match up the colors :3

OpenStudy (anonymous):

oh ok lol ya hold on

OpenStudy (anonymous):

y - 4 = 6(x - 7)

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

ok next what

zepdrix (zepdrix):

That's it :o the problem is finished.

OpenStudy (anonymous):

oh ok lol i have a 4 more could u help?

zepdrix (zepdrix):

ya i could spare a few i think :3

OpenStudy (anonymous):

do u want me to open a new question?

OpenStudy (anonymous):

Identify the point (x1, y1) from the equation: y + 8 = 3(x – 2) (Points : 1) (–2, 8) (8, –2) (–8,2) (2, –8)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

This is just like the last one, except in reverse. It'll be a little tricky since they threw a plus sign at us though.

zepdrix (zepdrix):

So remember that our standard form is:\[\Large\rm y-\color{orangered}{y_1}=\color{green}{m}(x-\color{royalblue}{x_1})\] See how I wrote MINUS y1? How would it change to positive? How would it change it addition?

OpenStudy (anonymous):

the x is 2 and the y is -8?

zepdrix (zepdrix):

Yes :) Good good good. The y would have to be negative!

OpenStudy (anonymous):

yay ok a couple more :)

OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (–2, 4); Slope: 5 (Points : 1) y + 4 = 5x – 10 y + 4 = 5(x + 2) y – 4 = 5(x + 2) y – 4 = 5(x – 2)

zepdrix (zepdrix):

lol they're all the same XD I'll bet you can do this one :3 what do you think?

OpenStudy (anonymous):

ok give me a sec. I'm asking so u can check :)

OpenStudy (anonymous):

y - 4 = 5(x - 2)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

This one is a little tricky,\[\Large\rm y-\color{orangered}{y_1}=\color{green}{m}(x-\color{royalblue}{x_1})\]\[\Large\rm y-\color{orangered}{4}=\color{green}{5}(x-\color{royalblue}{-2})\]So what is happening to the 2 by our x?

OpenStudy (anonymous):

IT BECOMES A PLUS SIGN

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

or no

zepdrix (zepdrix):

yes :) so that's the only thing we need to fix.

OpenStudy (anonymous):

oh ok :)

OpenStudy (anonymous):

Which equation is graphed here? algebra unit 7 lesson 8 assessment question 5 (Points : 1) y – 4 = (x + 2) y – 2 = (x + 4) y – 4 = (x + 2)

OpenStudy (anonymous):

thats the graph^

zepdrix (zepdrix):

Hmm none of your options have a slope value showing. Did they not paste correctly?